OpenStudy (anonymous):

A house painting company charges $376 plus $12 per hour. Another company $280 plus $15 per hour. A. how long is a job for which both companies will charge the same amount? B.What will the cost be?

OpenStudy (amistre64):

A = 376 + 12(h) B = 280 + 15(h) when does A = B? A = 376 + 12(h) = B = 280 + 15(h) 376 + 12(h) = 280 + 15(h), and solve for h